A person is ordered by doctor to take 15 units of vitamin A, 11 units of vitamin D, and 9 units of vitamin E each day. There are three brands of vitamin supplement available: brand X contains 2, 1, and 2 units, brand Y contains 3, 2, and 2 units and brand Z contains 4, 4, and 1 units of vitamin A, D, and E, respectively. How many pills of each brand will provide the required amounts of vitamin? ___ pills of brand X, ___ pills of brand Y and ___ pills of brand Z.

Answer by TimothyLamb(4379)

x = pills of brand X
y = pills of brand Y
z = pills of brand Z
---
units of A, D, and E per brand:
X: 2, 1, and 2
Y: 3, 2, and 2
Z: 4, 4, and 1
---
units of A, D, and E required by doctor:
15, 11, and 9
---
the linear system
2x + 3y + 4z = 15
1x + 2y + 4z = 11
2x + 2y + 1z = 9
---
put the system of linear equations into standard form
---
2x + 3y + 4z = 15
1x + 2y + 4z = 11
2x + 2y + 1z = 9
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 1
y = 3
z = 1
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---